An urn contains 5 black and 5 white balls. The probability of drawing two balls of the same colour
- A. 2/9
- B. 4/9
- C. 1/9
- D. 5/9
Answer: B. 4/9
Explanation: P(same colour) = P(2 Black) + P(2 White) = (5/10 × 4/9) + (5/10 × 4/9) = 20/90 + 20/90 = 40/90 = 4/9.